### v1

# RELATE: Relation Extraction in Biomedical Abstracts with LLMs and Ontology Constraints

Abstract
Biomedical knowledge graphs (KGs) are vital for drug discovery and clinical decision support but remain incomplete. Large language models (LLMs) excel at extracting biomedical relations, yet their outputs lack standardization and alignment with ontologies, limiting KG integration. We introduce RELATE, a three-stage pipeline that maps LLM-extracted relations to standardized ontology predicates using ChemProt and the Biolink Model. The pipeline includes: (1) ontology preprocessing with predicate embeddings, (2) similarity-based retrieval enhanced with SapBERT, and (3) LLM-based reranking with explicit negation handling. This approach transforms relation extraction from free-text outputs to structured, ontology-constrained representations. On the ChemProt benchmark, RELATE achieves 52% exact match and 94% accuracy@10, and in 2,400 HEAL Project abstracts, it effectively rejects irrelevant associations (0.4%) and identifies negated assertions. RELATE captures nuanced biomedical relationships while ensuring quality for KG augmentation. By combining vector search with contextual LLM reasoning, RELATE provides a scalable, semantically accurate framework for converting unstructured biomedical literature into standardized KGs.
keywords:
Relation Extraction, Large Language Models, Ontology, Biolink Protocol, Biomedical Knowledge Graphs
Data and Code Availability
Data used in this project are publicly available ChemProt dataset [^1] https://huggingface.co/datasets/bigbio/chemprot and PubMed abstracts available through the NIH Helping to End Addiction Long-term (HEAL) Initiative [^2] https://heal.nih.gov/. The code is available in a public Git repository [^3] https://github.com/RENCI-NER/pred-mapping/tree/multi-ontology.
Institutional Review Board (IRB)
This research analyzes public biomedical publication texts on PubMed and does not require IRB approval.
1 Introduction
Knowledge graphs (KGs) serve as structured repositories of facts, capturing entities and their relationships to support applications such as semantic search, recommendation systems, and drug repurposing (Olasunkanmi et al., 2024; Bizon et al., 2019). However, real-world KGs are often incomplete, with missing facts that limit their utility.
Unstructured biomedical text offers a vast, underutilized source for KG completion. Natural language processing techniques, particularly relation extraction (RE), play a central role in mining relationships between entities mentioned in the text. While Named Entity Recognition (NER) identifies entities (e.g., drugs, proteins, diseases), RE captures the predicates describing their relationships (e.g., treats, inhibits), which provide structure and meaning to extracted knowledge. Predicates thus connect entities into an interconnected biomedical KGs. Biomedical texts, however, are challenging to analyze due to domain-specific jargon, abbreviations, and subtle contextual nuances.
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in biomedical text understanding (Jahan et al., 2024). Nevertheless, current LLM-based RE methods often lack integration with domain ontologies, leading to high false-positive rates and inconsistent relationship representations. Standardized ontology frameworks such as the Biolink Model (Unni et al., 2022) define well-structured predicates for biomedical relationships (Bizon et al., 2019; Joachimiak et al., 2021; Reese et al., 2021; all, 2022). Yet, mapping free-text relations to ontology predicates remains a difficult task due to the semantic gap between natural language expressions and formal ontological representations.
To address these challenges, we propose RELATE, a three-stage pipeline that maps free-form biomedical relations to standardized ontology predicates. The key contributions of this work are:
  1. 1.
    Ontology-constrained relation standardization: RELATE introduces the first systematic framework for mapping LLM-extracted biomedical relations to established ontologies (ChemProt and Biolink), moving beyond relation classification toward ontology-grounded standardization. This design ensures semantic interoperability with biomedical KGs. While ChemProt has been used for supervised relation classification (Peng et al., 2019; Warikoo et al., 2021), and Biolink for knowledge graph schema definition (Joachimiak et al., 2021; Reese et al., 2021), our work represents the first application of these ontologies as standardization targets for free-text relation extraction, bridging the gap between unstructured text mining and structured knowledge representation.
  2. 2.
    Hybrid retrieval and efficient refinement: We develop a three-stage pipeline that combines ontology-driven preprocessing, SapBERT-enhanced hybrid vector retrieval, and LLM-based contextual reranking. RELATE leverages SapBERT for relation semantics for the first time, and applies computationally intensive reasoning only to small candidate sets ($k=10$), achieving both scalability and accuracy.
  3. 3.
    Comprehensive evaluation: We conduct dual-ontology assessment on ChemProt (Peng et al., 2019) benchmark and 2,400 real-world HEAL Project abstracts. Integration with an established biomedical KG (ROBOKOP) is also planned to further demonstrate its translational impact.
[…]
fig:workflow Refer to caption
Figure 1: The three-stage RELATE pipeline: (1) ontology preprocessing, (2) similarity-based retrieval, and (3) contextual refinement. Ontology preprocessing generates embeddings for predicates and their negated variants. It updates only if the ontology schema or embedding models change. Given an input triple with abstract context, RELATE performs similarity-based retrieval by embedding the relation and retrieving top-$k$ ontology candidates. In the final stage, contextual refinement reranks these candidates with an LLM using the full abstract context, producing the most semantically appropriate ontology predicate.
2 Related Work
No existing framework systematically bridges free-text biomedical RE with standardized ontological predicates. While zero-shot LLMs (Li et al., 2025) can outperform fine-tuned models on smaller datasets, they struggle when confronted with complex, domain-specific tasks that demand specialized knowledge (Chen et al., 2024a; Jahan et al., 2024). Further research has explored adaptive instruction-rich prompting (Zhou et al., 2024; Tao et al., 2024) and soft prompt-based learning approaches that allow models to automatically optimize prompts rather than relying on hand-crafted instructions (Peng et al., 2024). These methods show promise but have not achieved systematic biomedical ontology integration.
Fine-tuning approaches have demonstrated strong performance on domain-specific benchmarks. Techniques such as adaptive document mapping, ensemble learning with attention mechanisms, and specialized training procedures (Shang et al., 2025; Zhong et al., 2024; Orlova and Orlov, 2025) improve classification accuracy. However, they require extensive retraining and do not address the core challenge of standardizing predicates extracted from free-text relations.
Recent work in LLM-based relation extraction has explored various validation and standardization approaches. RelCheck (Ourekouch et al., 2025) addresses low-confidence predictions from pretrained models using ontology-guided LLM validation, though it focuses on confidence rather than standardization. SPIREX (Cavalleri et al., 2024) applies schema-constrained prompts with graph machine learning validation but restricts itself to RNA-specific domains. Graph-augmented approaches such as RAG-enhanced LLMs for automated ontology extension (Georgakopoulos et al., 2025); structure-oriented RAG for KG-LLM co-learning (Yang et al., 2024); and integrated GNNs with LLM-generated support documents (Dong et al., 2024), and LLM-driven ontology enrichment pipelines (Kollapally et al., 2025).
However, these approaches primarily target domain-specific extraction, confidence validation, or knowledge system extension rather than systematic standardization of extracted relationships to established ontological frameworks. Multi-channel neural architectures combine textual and graph structures but focus primarily on structural features rather than systematic predicate mapping. Caufield et al. (2024) applied structured prompt interrogation for ontology-grounded annotations from scientific publications, yet their focus remained on entity linking rather than comprehensive relation standardization. Even enhancements to systems such as SemRep (Ming et al., 2024) that leverage contrastive learning operate within existing predicate frameworks rather than enabling mapping to broader ontological schemas.
3 RELATE Methodology
Biomedical relation extraction (RE) systems face a critical challenge: while LLMs excel at identifying subject–object relationships in scientific text, the extracted relations often lack standardization and fail to align with biomedical ontologies. This disconnect limits their value for KG construction and augmentation.
To address this limitation, we propose a three-stage pipeline (Figure LABEL:fig:workflow) that bridges free-text RE with standardized predicates through an ontology-driven protocol. The pipeline consists of: (1) Ontology Preprocessing, which generates positive and negative descriptors for predicates and embeds them into a searchable space; (2) Similarity-Based Retrieval, which efficiently narrows candidate predicates using vector search enhanced with biomedical embeddings; and (3) Contextual Refinement, which leverages LLM reasoning and abstract-level context to rerank candidates and select the most appropriate predicate.
3.1 Stage 1: Ontology Preprocessing
The first stage of our pipeline (Alg. 1) extracts predicates with positive and negative descriptors to encode domain knowledge, and transforms them into a searchable embedding database suitable for relation mapping.
We begin by collecting all predicate definitions from biomedical ontologies such as the Biolink Model toolkit (Unni et al., 2022) or ChemProt (Krallinger et al., 2017), thereby constructing a base vocabulary of biomedical relationships. We will denote the set of all ontology predicates as $\mathcal{R}$. Each predicate $r\in\mathcal{R}$ represents a distinct biological or medical relationship (e.g., affects, inhibits) that can exist between entities in the KG.
The descriptor dataset $\mathcal{D}_{r}^{+}$ is constructed by gathering existing textual descriptions for each predicate $r$. For example, one descriptor for the predicate affects may be “describes an entity that has an effect on the state or quality of another existing entity”. These descriptions are sourced from multiple authoritative biomedical sources. In the Biolink example, this includes official documentation and standardized vocabularies, while in the ChemProt case, online relation category descriptions suffice. The distributions of the predicate descriptors are presented in Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal. These aggregated descriptions form the positive dataset $\mathcal{D}_{r}^{+}$ and capture the semantic diversity with which biomedical relationships appear in scientific literature.
To complement the positive dataset, we introduce a novel strategy for generating negative descriptors. For each descriptor in $D_{r}^{+}$, an LLM produces natural, semantically coherent negations using a structured prompt (Figure LABEL:fig:negation_prompt). For example, the earlier descriptor becomes “describes an entity that does not have an effect on the state or quality of another existing entity.” These negated descriptors are stored in a separate dictionary $D_{r}^{-}$ to serve as contrastive examples that help the embedding model distinguish valid from invalid predicate mappings. $D_{r}^{-}$ is labeled by $r$ + “_NEG” (i.e. affects $\to$ affects_NEG) to maintain clear separation from positive instances.
Finally, every descriptor in $D^{+}_{r}$ and $D^{-}_{r}$ is transformed into $d$-dimensional vector representations using the embedding model $f_{\theta}$. Instead of aggregating these descriptor embeddings to a single predicate embedding, we store all descriptor embeddings, $\mathcal{V}_{r}$, in a searchable database for the next stage. To address limitations of general-purpose embedding models (e.g., nomic-text-embed or bge-m3) in biomedical domains, we introduce an (optional) hybrid retrieval enhancement based on SapBERT (Liu et al., 2021), a model specialized in biomedical entity linking. In this setup (Figure LABEL:fig:workflowplus), SapBERT was fine-tuned on $D^{+}_{r}$ and $D^{-}_{r}$ to generate a second set of embeddings $\mathcal{V}^{\prime}_{r}$. While SapBERT has demonstrated effectiveness in biomedical entity alignment (Liu et al., 2021), this work presents its first application to relation extraction through predicate embedding enhancement. Our hybrid approach leverages SapBERT’s biomedical domain knowledge to improve relation semantic similarity beyond general-purpose embedding models.
[…]
fig:negation_prompt You are a biomedical researcher extracting negations of ontological predicates.
Your Task:
Given a description, return its natural negation.
Rules:
1. Preserve the meaning but negate the entire description.
3. Do not summarize or change the structure of the descriptor text.
4. If there is not enough information to create a negation, your response should be "NOT ENOUGH INFORMATION"
4. Only return the negation—no explanations or extra text.
Examples:
- "has effect" → "does not have effect"
- "during which ends" → "during which does not ends"
- "happens during" → "does not happen during"
Input: "{descriptor_text}"
Output: A JSON object with these exact keys and format:
{{"negation_of_the_descriptor_text": "negated version" or "NOT ENOUGH INFORMATION"}}
Figure 2: Negation generation prompt used in Stage 3.1 for creating negative descriptors from positive descriptors.
3.2 Stage 2: Similarity-Based Retrieval
For every predicate $r\in\mathcal{R}$, Stage 1 has produced positive and negative descriptor dictionaries $\mathcal{D}_{r}^{+}$, $\mathcal{D}_{r}^{-}$ and corresponding descriptor embedding vectors $\mathcal{V}_{r}$. We now leverage these data to retrieve candidate ontology predicates for incoming extracted relations.
The extracted relations come in a quadruple $(s,o,T,a)$: $s$ subject entity, $o$ object entity, $T$ the free-form relation text, and $a$ the abstract context from which the relation was identified. For example, $T$ could be “increases level of,” which does not exist in Biolink ontology in this format. First, free-form relation text $T$ is embedded to query vector $q=f_{\theta}(T)$. We use the same model $f_{\theta}$ from Stage 1, ensuring semantic consistency between query relations and stored ontology predicates.
Once embedded, the query vector $q$ is compared against all predicate descriptor embeddings in $\mathcal{V}_{r},\;~\forall r\in\mathcal{R}$. This similarity search retrieves the top-$k$ candidates (typically $k=10$) using cosine similarity. If multiple descriptors of the same predicate are in the top-$k$ results, they are collapsed into a single predicate candidate. When SapBERT enhancement is on, the top-$k$ candidates in $\mathcal{V}^{\prime}_{r},\;~\forall r\in\mathcal{R}$ are retrieved in parallel. Then, the two sets of candidates are merged and de-duplicated to identify (up to) 2kk candidate predicates.
By narrowing the search space from hundreds of ontology predicates to a small candidate set, this stage provides an efficient yet semantically informed mechanism for predicate selection. At this point, however, the search operates without the benefit of contextual information from the original biomedical abstract, focusing only on lexical and semantic similarity.
3.3 Stage 3: Contextual Refinement
[…]
fig:reranking_prompt You are an expert in biomedical relationships. Based on the text below:
Subject: {subject}
Object: {object}
Original Relationship: {relationship}
Abstract: {abstract}
Candidate Predicates:
{choices_str}
Instructions:
- Choose the best predicate from the list that matches the intended meaning and direction.
- If the original relationship implies negation (e.g., "does not cause"), select the matching base predicate, but set "negated" to "True".
- If no match exists, return ‘"mapped_predicate": "none"‘.
Respond with ONLY this JSON object:
{{"mapped_predicate": "one of the predicate keys or ’none’", "negated": "True" or "False"}}
Figure 3: Contextual reranking prompt used in Stage 3 for LLM-based predicate selection with explicit negation handling.
The final stage addresses the key limitation of similarity-based retrieval: its reliance on context-free embeddings. While Stage 2 efficiently narrows the predicate space to a small candidate set, it does not incorporate the biomedical context in which the relation was expressed. Stage 3 introduces contextual refinement, where an LLM evaluates and re-ranks the top-$k$ candidates against the full relation context.
The LLM is given structured prompt (Figure LABEL:fig:reranking_prompt) along with the following information: the top candidate predicates retrieved by similarity-based search and the full RE quadruple $(s,o,T,a)$ including the abstract context. By explicitly grounding the candidate predicates in this richer context, the LLM can better assess which ontology predicate is most appropriate for the specific biomedical scenario.
The LLM then performs re-ranking of the candidate set, selecting the single predicate that best aligns with the domain context. Importantly, the system includes a NONE option, allowing the model to reject all candidates if none provide a semantically valid match. This safeguard prevents the forced assignment of predicates in cases where the extracted relation is either too domain-specific, erroneous, or incompatible with the ontology.
The output of Stage 3 is a contextually validated predicate $r^{*}$ that integrates both semantic similarity and biomedical context. This refinement ensures that only relationships consistent with both the ontology and the source text are admitted into the knowledge graph, thereby improving accuracy and reducing false positives.
4 Experiments
[…]
tab:chemprot_overall LLM Embedding Exact Match a@1 a@3 a@5 a@10 MRR MedGemma 27B nomic 0.464 0.528 0.686 0.744 0.755 0.612 MedGemma 27B bge-m3 0.467 0.468 0.686 0.788 0.812 0.590 MedGemma 27B nomic + SapBERT 0.520 0.100 0.328 0.688 0.914 0.300 MedGemma 27B bge + SapBERT 0.516 0.100 0.321 0.691 0.940 0.302
Table 1: RELATE performance on ChemProt. Exact Match: accuracy of RELATE; a@$k$: accuracy in top-$k$ candidates before LLM-based refinement in Stage 3 of RELATE; MRR: Mean reciprocal rank.
We assess RELATE on two independent datasets with different ontology models, as well as four different embedding model configurations.
4.1 Datasets
The ChemProt dataset (Krallinger et al., 2017) provides annotated chemical-protein relations representing diverse pharmacological and biochemical relationship types, including regulators, modulators, agonists, antagonists, and substrate interactions. We treat the information in ChemProt as ground truth abstract-predicate pairs for evaluation purposes. Standard metrics in information retrieval are used, such as exact match accuracy, accuracy@$k$, and Mean Reciprocal Rank (MRR). Recall from Figure LABEL:fig:reranking_prompt that Stage 3 of RELATE only outputs its top choice predicate $r^{*}$ or NONE, and does not return a fully re-ranked or re-ordered list of $k$ candidates. Therefore, exact match accuracy is calculated with respect to the final output of RELATE, $r^{*}$, but accuracy@$k$ and MRR are calculated using the top-$k$ candidates after Stage 2, before LLM-based Stage 3. This will allow us to analyze the effect of LLM-based refinement in isolation as well.
Next, we apply RELATE to 2,400 PubMed abstracts from the HEAL Project [^4] https://heal.nih.gov targeting opioid use disorder research. This real-world dataset reflects the practical application scenario where relationships extracted from literature must be mapped to standardized Biolink predicates for integration into a large biomedical KG [^5] Name hidden during double blind review process.. Unlike ChemProt, the HEAL abstracts lack ground-truth annotations, requiring qualitative assessment of RELATE’s outputs.
Importantly, this corpus captures the complexity and noisiness of real biomedical literature. In contrast to benchmark datasets, these abstracts often contain methodological statements or negative scientific findings (e.g., drug does not affect condition, gene not associated with disease). Such cases underscore the necessity of RELATE’s rejection and negation-handling capabilities to ensure only semantically appropriate biomedical relationships enter the KG. As part of the integration plan, the abstract authors and Subject Matter Experts (SMEs) are being recruited to systematically evaluate the quality and reliability of the extracted triples.
4.2 LLM and Embedding Models
We experimented with MedGemma (4B), LLaMA2 (7B), LLaMA3 (8B), LLaMA3 (70B), MedGemma (27B), and OpenAI GPT-4o-mini (unknown number of parameters). The last two demonstrated greater consistency and fewer hallucinations compared to the other models. Since those two models had very similar performances, we only report results using MedGemma in this work. However, note that any other LLM could be substituted to reproduce the pipeline. MedGemma-27B is Google’s medical language model (Sellergren et al., 2025), trained exclusively on medical text and optimized for efficient inference through architectural improvements in performance and computational efficiency.
As with the LLMs, we also experimented with several embedding approaches. While we prioritize accessibility through ollama-based open-access models, we tested OpenAI’s text-embed-large models and others with 8,192-token context windows that provide general-purpose semantic understanding. These include nomic-embed-text (Zach et al., ), a 768-dimensional embedding model, and bge-m3, a 1,024-dimensional multilingual embedding model for diverse textual contexts (Chen et al., 2024b). Since nomic-embed-text and bge-m3 achieved comparable performance, we report only the results obtained with the 768-dimensional nomic-embed-text in this work.
For (optional) hybrid retrieval, we experimented with adding fine-tuned SapBERT (Liu et al., 2021). We initialized the model with BiomedNLP-PubMedBERT-base-uncased-abstract -fulltext and optimized it using a contrastive learning framework. Specifically, we employed multi-similarity loss with hard negative mining, which encourages the model to better distinguish between semantically close but incorrect pairs. To keep training efficient while still capturing relevant semantic features, we limited input sequences to 25 tokens and represented each instance using the [CLS] embedding.
Training was carried out for 10 epochs with a batch size of 256 and a learning rate of $2\times 10^{-5}$. We used mixed-precision training to reduce computational overhead and set a fixed random seed for reproducibility.
5 Results and Discussions
LLM-based Refinement is Necessary.
Table LABEL:tab:chemprot_overall illustrates how embedding choice shapes RELATE’s performance on the ChemProt benchmark. Baseline configurations with nomic and bge-m3 produce stronger top-rank accuracy (a@1 up to 0.528, MRR $\approx$ 0.6). In contrast, SapBERT-enhanced embeddings yield slightly higher Exact Match (0.520) and dramatically improve retrieval coverage, with a@10 reaching 0.940. This indicates that SapBERT ensures the correct ontology predicate is almost always retrieved, though often ranked lower in the list (a@1 drops to 0.1, MRR $\approx$ 0.3).
These results highlight a trade-off: baseline embeddings rank the correct answer more accurately at the top, while SapBERT embeddings broaden coverage and guarantee inclusion within candidate sets. RELATE resolves this through its LLM-based refinement stage, which reranks SapBERT’s expanded candidate lists in context. This allows RELATE to recover early precision while preserving high recall, transforming broad retrieval into precise ontology-constrained mappings. In short, SapBERT enhances retrieval coverage, but RELATE’s LLM-based refinement is essential to convert this coverage into accurate, ontology-aligned relations.
When Does RELATE Reject Candidates?
RELATE’s rejection capability (NONE) represents a critical quality control mechanism. Out of 2,400 PubMed abstracts from the HEAL project, 10 relationships were appropriately rejected as unsuitable for Biolink predicate mapping (rejection rate:  0.4%). Analysis of these rejections reveals contextual understanding that distinguishes between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
The rejected relationships fell into distinct categories reflecting non-biomedical associations: genetic nomenclature specifications (3 cases), clinical procedures and documentation (4 cases, including “Clinicians prescribe Opioid” and “Women on BUP had documented Opioid misuse”), methodological relationships (2 cases, such as “Anisomycin used to detect GNPTG”), and specialized reporting contexts (1 case). These rejections demonstrate the pipeline’s ability to recognize when extracted relationships represent procedural, methodological, or administrative associations rather than biological or medical relationships suitable for KG representation.
These rejections occurred despite high embedding similarity scores, demonstrating the reranker’s comprehensive contextual analysis. The low rejection rate (0.4%) indicates that the majority of extracted relationships represent valid biomedical associations while maintaining strict quality standards for edge cases requiring specialized handling.
Chemprot analysis also reveals potential false rejections where high similarity scores failed to produce mappings. Cases like “loperamide modifies corticotropin-releasing hormone actio“ (0.661 score for “modulator“) and “ammonium sulfate reduces CBG concentrations“ (0.632 for “downregulator“) were rejected despite appropriate candidates. While some discrepancies reflect stricter semantic criteria, others suggest overly conservative contextual reranking, highlighting evaluation challenges with traditional benchmarks.
When Does RELATE Choose Negated Candidates?
Out of 2,400 PubMed abstracts from the HEAL project, 77 relationships were correctly flagged as negated assertions, revealing how well the system processes the nuanced ways of reporting biomedical negative findings. Most prevalent was “not associated with”, which accounted for nearly half of the negated cases (32 instances). Beyond these straightforward cases, we found the pipeline successfully navigating more complex terrain. Other negations like “does not affect” appeared in 8 cases, while direct causal negations (“does not cause”) showed up 7 times. The remaining quarter of cases presented genuine linguistic challenges. Consider phrases like “remained impaired despite treatment” or “not necessary for acute insulin-induced uptake”. These require contextual understanding that goes well beyond pattern matching. The system had to recognize that “fails to affect” and “unable to deliver” carry negative semantic weight, even when the word “not” never appears.
Perhaps most importantly, RELATE maintained conceptual accuracy throughout this process. When it encountered “Drug X does not treat Disease Y,” it correctly identified the underlying relationship as therapeutic (mapping to treats) while flagging the negation. This dual parsing of both the relationship type and its polarity proved consistent across our dataset. The most frequent base relationships were association patterns (28 cases) and effect relationships (18 cases), followed by causal relationships (8 cases). The other application capability of negation handling includes findings on use, such as which treatments fail, which genetic variants are not linked to diseases, and which interventions prove ineffective.
SAPBERT Enhancement Helps.
To evaluate SapBERT’s effect, we compared 2,139 predicate assignments between bge-m3 and bge-m3 + SapBERT. Their disagreement patterns reveal four key improvements by SAPBERT aligned with biomedical expert judgment. Clinical terminology refinement was most prominent, with 28 cases shifting from associated with: increased likelihood of to predisposes to condition. Domain specificity enhancement appeared in 12 cases where generic associated with became gene associated with condition for relationships like “DRD4 associated with Heroin Addiction,” recognizing genetic contexts require more informative predicates.
Evidence-based caution moderated 12 strong causal claims, changing causes to contributes to for relationships like “MDMA causes Depression,” reflecting the complex multifactorial causation typical in biomedicine, where direct causality claims often overstate evidence strength. Mechanistic precision enhanced molecular relationship descriptions, with 10 cases shifting from affects: increased abundance to affects: increased activity or abundance for relationships like “Betaine enhances phosphorylation of STAT3,” capturing both quantitative and functional effects in biochemical interactions. These systematic patterns demonstrate that SapBERT’s biomedical entity embeddings guide predicate selection toward clinically appropriate, evidence-aware terminology that better reflects expert knowledge of biomedical relationship complexity.
6 Conclusion
We introduced RELATE, a three-stage pipeline for automated ontology-constrained predicate mapping in biomedical relation extraction. By combining ontology-driven preprocessing, efficient vector search, and context-aware LLM reranking, RELATE systematically converts free-text relations into ontology-aligned, interoperable knowledge graph edges. Experiments on ChemProt demonstrate the trade-offs between early precision and retrieval depth, while application to 2,400 HEAL Project abstracts illustrates RELATE’s ability to handle noisy real-world biomedical literature, including identification of negated assertions and appropriate rejection of non-biomedical associations. Together, these results establish a foundation for scalable, ontology-constrained relation extraction and highlight a path forward for enriching biomedical knowledge graphs with high-quality, standardized information. Future work includes improving computational overhead, designing a scalable evaluation that combines expert judgment with automated consistency checks, and integration into the existing large biomedical KG (ROBOKOP).
[…]
We thank the LitCoin project partners for their immense contributions, the HEAL initiative consortium for providing access to opioid research abstracts. Also, we acknowledge the RENCI’s Data Analytics group and Knowledge Graph Working Group for constructive reviews. This research was supported by NIH #75N95023C00032 (LITCOIN)
Appendix A Relation Extraction Pipeline Stages Algorithms
Input: Biolink Model YAML $\mathcal{B}$, Biolink Toolkit $\mathcal{T}$, LLM $\mathcal{M}$, Embedding model $f_{\theta}$
Output: Embedding store $\mathcal{V}$, Positive descriptors $\mathcal{D}^{+}$, Negative descriptors $\mathcal{D}^{-}$
[…]
$\mathcal{D}^{+}\leftarrow\emptyset$;
for each predicate mapping $r\in\mathcal{R}$ do
 $\text{predicate}\leftarrow r.\text{predicate}$;
 $\text{descriptions}\leftarrow$ GatherDescriptions(predicate)($\text${predicate});
 $\mathcal{D}^{+}[\text{predicate}]\leftarrow\text{descriptions}$;
end for
$\mathcal{D}^{-}\leftarrow\emptyset$;
for each predicate $p\in\mathcal{D}^{+}$ do
 $\text{pos\_descriptions}\leftarrow\mathcal{D}^{+}[p]$;
 $\text{neg\_descriptions}\leftarrow\emptyset$;
 for each description $d\in\text{pos\_descriptions}$ do
    $\text{prompt}\leftarrow$ CreateNegationPrompt(d)(d);
    $d^{neg}\leftarrow\mathcal{M}(\text{prompt})$;
    $\text{neg\_descriptions}\leftarrow\text{neg\_descriptions}\cup\{d^{neg}\}$;
   end for
 $\text{neg\_key}\leftarrow p+\text{"\_NEG"}$;
 $\mathcal{D}^{-}[\text{neg\_key}]\leftarrow\text{neg\_descriptions}$;
end for
$\mathcal{V}\leftarrow\emptyset$;
for each predicate $p\in\mathcal{D}^{+}\cup\mathcal{D}^{-}$ do
 $\text{descriptions}\leftarrow\mathcal{D}^{+}[p]\text{ if }p\in\mathcal{D}^{+}\text{ else }\mathcal{D}^{-}[p]$;
 $\text{combined\_text}\leftarrow$ CombineDescriptions(descriptions)($\text${descriptions});
 $\mathbf{v}_{p}\leftarrow f_{\theta}(\text{combined\_text})$;
 $\mathcal{V}\leftarrow\mathcal{V}\cup\{(\text{label}:p,\text{embedding}:\mathbf{v}_{p})\}$;
end for
return $\mathcal{V},\mathcal{D}^{+},\mathcal{D}^{-}$;
Algorithm 1 Ontology Preprocessing
Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Predicate embeddings $\mathcal{V}$, Embedding model $f_{\theta}$, Top-k parameter $k$
Output: $\mathcal{C}$
// Relationship Embedding Phase
$\mathcal{R}_{embed}\leftarrow\emptyset$;
for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
 $\mathbf{v}_{r_{i}}\leftarrow f_{\theta}(r_{i})$;
 $\mathcal{R}_{embed}\leftarrow\mathcal{R}_{embed}\cup\{(i,\mathbf{v}_{r_{i}})\}$;
end for
// Cosine Similarity Search and Top-k Candidate Retrieval
$\mathcal{C}\leftarrow\emptyset$;
for each $(i,\mathbf{v}_{r_{i}})\in\mathcal{R}_{embed}$ do
 $\text{similarities}\leftarrow\emptyset$;
 for each $(\text{label}_{j},\mathbf{v}_{j})\in\mathcal{V}$ do
    $\text{sim}_{j}\leftarrow\cos(\mathbf{v}_{r_{i}},\mathbf{v}_{j})=\frac{\mathbf{v}_{r_{i}}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{r_{i}}\|\|\mathbf{v}_{j}\|}$;
    $\text{similarities}\leftarrow\text{similarities}\cup\{(\text{label}_{j},\text{sim}_{j})\}$;
   end for
 $\text{sorted}\leftarrow$ SortDescending(similarities)($\text${similarities});
 $\text{candidates}_{i}\leftarrow\text{sorted}[1:k]$;
 $\mathcal{C}\leftarrow\mathcal{C}\cup\{(i,\text{candidates}_{i})\}$;
end for
return $\mathcal{C}$;
Algorithm 2 Similarity-Based Retrieval
Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Candidate sets $\mathcal{C}$, LLM $\mathcal{M}$
Output: Mapped triples $\mathcal{T}^{*}=\{(s_{i},o_{i},r_{i}^{*},a_{i})\}_{i=1}^{n}$
$\mathcal{T}^{*}\leftarrow\emptyset$;
for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
 $\text{candidates}_{i}\leftarrow\mathcal{C}[i]$;
[…]
 // Prompt Construction
 $\text{prompt}\leftarrow$ ”Given the biomedical relationship context:”;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Subject: "}+s_{i}+\text{" Object: "}+o_{i}$;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Relationship: "}+r_{i}+\text{" Abstract: "}+a_{i}$;
 $\text{prompt}\leftarrow\text{prompt}+\text{" Select best Biolink predicate or NONE:"}$;
 for $j=1$ to $|\text{predicate\_list}|$ do
    $\text{prompt}\leftarrow\text{prompt}+j+\text{". "}+\text{predicate\_list}[j]$;
   end for
 $\text{prompt}\leftarrow\text{prompt}+(|\text{predicate\_list}|+1)+\text{". NONE"}$;
 // LLM Selection and Validation
 $\text{response}\leftarrow\mathcal{M}(\text{prompt})$;
[…]
 if $r_{i}^{*}=\text{"NONE"}$ or $r_{i}^{*}\notin\text{predicate\_list}$ then
    $r_{i}^{*}\leftarrow\emptyset$;
   end if
 $\mathcal{T}^{*}\leftarrow\mathcal{T}^{*}\cup\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
end for
return $\mathcal{T}^{*}$;
Algorithm 3 Contextual Refinement
Appendix B SAPBERT Workflow Enhancement
[…]
fig:workflowplus Refer to caption
Figure 4: SapBERT Augmented System workflow diagram showing the three-stage pipeline as in Figure LABEL:fig:workflow: (1) Ontology Preprocessing, (2) Similarity-Based Retrieval, and (3) Contextual Refinement. Just as in Figure LABEL:fig:workflow, stage 3.1 changes only if the ontology schema or the LLM and embedding Model changes. However, the one-stage preprocessing of raw ontology predicate data involves two embedding generations- LLM-generated embedding and SapBERT finetuned stored embedding. Also, given a raw input triple with abstract context, the frequently executed stage—similarity-based retrieval and contextual refinement—encompasses the input relation’s dual embedding transformation, dual similarity search on the pair of stored embeddings, merger of top-k candidates from each search, and final contextual reranking.
Appendix C Predicate and Descriptor Distribution
[…]
tab:top_predicates_distribution_chemprot Predicate Descriptors Percentage (%) substrate 6 20.00 agonist 5 16.67 upregulator 4 13.33 downregulator 4 13.33 regulator 3 10.00 part of 2 6.67 antagonist 2 6.67 modulator 2 6.67 cofactor 2 6.67
Table 2: Chemprot Dataset Top Predicates by Descriptor Count and Percentage Contribution
[…]
tab:top_predicates_distribution_heal Predicate Descriptors Percentage (%) coexists with 178 39.04 related to 96 21.05 located in 40 8.77 part of 34 7.46 temporally related to 30 6.58 has part 21 4.61 precedes 15 3.29 causes 15 3.29 affects 14 3.07 has output 13 2.85
Table 3: Biolink Dataset Top Predicates by Descriptor Count and Percentage Contribution
Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal showed the top 10 predicates that account for the majority of descriptors in the dataset. substrate and agonist together accounting for over one-third of all descriptors in Table LABEL:tab:top_predicates_distribution_chemprot while Table LABEL:tab:top_predicates_distribution_heal dominated by predicates such as coexists with and related to.

### v2

@@
 Biomedical knowledge graphs (KGs) are vital for drug discovery and clinical decision support but remain incomplete.
-Large language models (LLMs) excel at extracting biomedical relations, yet their outputs lack standardization and alignment with ontologies, limiting KG integration.
-We introduce RELATE, a three-stage pipeline that maps LLM-extracted relations to standardized ontology predicates using ChemProt and the Biolink Model.
+Large language models (LLMs) excel at extracting biomedical relations, yet their outputs lack standardization and alignment with ontologies, limiting KG integration with free texts.
+We introduce RELATE, a three-stage pipeline that maps LLM-extracted relations to standardized ontology predicates, e.g., the Biolink Model.
 The pipeline includes: (1) ontology preprocessing with predicate embeddings, (2) similarity-based retrieval enhanced with SapBERT, and (3) LLM-based reranking with explicit negation handling.
-This approach transforms relation extraction from free-text outputs to structured, ontology-constrained representations.
+This approach performs relation extraction from free-text outputs to structured, ontology-constrained representations.
 On the ChemProt benchmark, RELATE achieves 52% exact match and 94% accuracy@10, and in 2,400 HEAL Project abstracts, it effectively rejects irrelevant associations (0.4%) and identifies negated assertions.
@@
 keywords:
-Relation Extraction, Large Language Models, Ontology, Biolink Protocol, Biomedical Knowledge Graphs
+Relation Extraction, Large Language Models, Ontology, Biolink Model, Biomedical Knowledge Graphs
 Data and Code Availability
 Data used in this project are publicly available ChemProt dataset [^1] https://huggingface.co/datasets/bigbio/chemprot and PubMed abstracts available through the NIH Helping to End Addiction Long-term (HEAL) Initiative [^2] https://heal.nih.gov/.
-The code is available in a public Git repository [^3] https://github.com/RENCI-NER/pred-mapping/tree/multi-ontology.
+The code is available in a public Git repository [^3] https://github.com/RENCI-NER/pred-mapping/releases/tag/v1.0.
 Institutional Review Board (IRB)
@@
 1 Introduction
-Knowledge graphs (KGs) serve as structured repositories of facts, capturing entities and their relationships to support applications such as semantic search, recommendation systems, and drug repurposing (Olasunkanmi et al., 2024; Bizon et al., 2019).
+Knowledge graphs (KGs) serve as structured repositories of facts, capturing entities and their relationships to support applications such as semantic search, recommendation systems, and drug repurposing (olasunkanmi2024explainable; bizon2019robokop).
 However, real-world KGs are often incomplete, with missing facts that limit their utility.
@@
 Biomedical texts, however, are challenging to analyze due to domain-specific jargon, abbreviations, and subtle contextual nuances.
-Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in biomedical text understanding (Jahan et al., 2024).
+Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in biomedical text understanding (jahan2024comprehensive).
 Nevertheless, current LLM-based RE methods often lack integration with domain ontologies, leading to high false-positive rates and inconsistent relationship representations.
-Standardized ontology frameworks such as the Biolink Model (Unni et al., 2022) define well-structured predicates for biomedical relationships (Bizon et al., 2019; Joachimiak et al., 2021; Reese et al., 2021; all, 2022).
+Standardized ontology frameworks such as the Biolink Model (unni2022biolink) define well-structured predicates for biomedical relationships (bizon2019robokop; joachimiak2021kg; reese2021kg; alliance2022harmonizing).
 Yet, mapping free-text relations to ontology predicates remains a difficult task due to the semantic gap between natural language expressions and formal ontological representations.
@@
 This design ensures semantic interoperability with biomedical KGs.
-While ChemProt has been used for supervised relation classification (Peng et al., 2019; Warikoo et al., 2021), and Biolink for knowledge graph schema definition (Joachimiak et al., 2021; Reese et al., 2021), our work represents the first application of these ontologies as standardization targets for free-text relation extraction, bridging the gap between unstructured text mining and structured knowledge representation.
+While ChemProt has been used for supervised relation classification (peng2019transfer; warikoo2021lbert), and Biolink for KG schema definition (joachimiak2021kg; reese2021kg), our work represents the first application of these ontologies as standardization targets for free-text relation extraction, bridging the gap between unstructured text mining and structured knowledge representation.
   2. 2.
@@
   3. 3.
-    Comprehensive evaluation: We conduct dual-ontology assessment on ChemProt (Peng et al., 2019) benchmark and 2,400 real-world HEAL Project abstracts.
-Integration with an established biomedical KG (ROBOKOP) is also planned to further demonstrate its translational impact.
+    Comprehensive evaluation: We conduct dual-ontology assessment on ChemProt (peng2019transfer) benchmark and 2,400 real-world HEAL Project abstracts.
+Integration with an established, large-scale biomedical KG (ROBOKOP) is also planned to further demonstrate its translational impact.
+These capabilities enable downstream applications in literature-based discovery, biomedical KG construction, and multi-source knowledge integration; see Appendix A.
 […]
@@
 It updates only if the ontology schema or embedding models change.
-Given an input triple with abstract context, RELATE performs similarity-based retrieval by embedding the relation and retrieving top-$k$ ontology candidates.
+Given an input quadruple—subject, object, relation text, and abstract context—RELATE performs similarity-based retrieval by embedding the relation and retrieving top-$k$ ontology candidates.
 In the final stage, contextual refinement reranks these candidates with an LLM using the full abstract context, producing the most semantically appropriate ontology predicate.
 2 Related Work
-No existing framework systematically bridges free-text biomedical RE with standardized ontological predicates.
-While zero-shot LLMs (Li et al., 2025) can outperform fine-tuned models on smaller datasets, they struggle when confronted with complex, domain-specific tasks that demand specialized knowledge (Chen et al., 2024a; Jahan et al., 2024).
-Further research has explored adaptive instruction-rich prompting (Zhou et al., 2024; Tao et al., 2024) and soft prompt-based learning approaches that allow models to automatically optimize prompts rather than relying on hand-crafted instructions (Peng et al., 2024).
+No existing framework systematically bridges free-text biomedical relation extraction (RE) with standardized ontological predicates.
+While zero-shot LLMs (li2025grag) can outperform fine-tuned models on smaller datasets, they struggle when confronted with complex, domain-specific tasks that demand specialized knowledge (chen2024benchmarking; jahan2024comprehensive).
+Further research has explored adaptive instruction-rich prompting (zhou2024leap; tao2024graphical) and soft prompt-based learning approaches that allow models to automatically optimize prompts rather than relying on hand-crafted instructions (peng2024model).
 These methods show promise but have not achieved systematic biomedical ontology integration.
 Fine-tuning approaches have demonstrated strong performance on domain-specific benchmarks.
-Techniques such as adaptive document mapping, ensemble learning with attention mechanisms, and specialized training procedures (Shang et al., 2025; Zhong et al., 2024; Orlova and Orlov, 2025) improve classification accuracy.
+Techniques such as adaptive document mapping, ensemble learning with attention mechanisms, and specialized training procedures (shang2025biomedical; zhong2024leveraging; orlova2025challenges) improve classification accuracy.
 However, they require extensive retraining and do not address the core challenge of standardizing predicates extracted from free-text relations.
-Recent work in LLM-based relation extraction has explored various validation and standardization approaches.
-RelCheck (Ourekouch et al., 2025) addresses low-confidence predictions from pretrained models using ontology-guided LLM validation, though it focuses on confidence rather than standardization.
-SPIREX (Cavalleri et al., 2024) applies schema-constrained prompts with graph machine learning validation but restricts itself to RNA-specific domains.
-Graph-augmented approaches such as RAG-enhanced LLMs for automated ontology extension (Georgakopoulos et al., 2025); structure-oriented RAG for KG-LLM co-learning (Yang et al., 2024); and integrated GNNs with LLM-generated support documents (Dong et al., 2024), and LLM-driven ontology enrichment pipelines (Kollapally et al., 2025).
-However, these approaches primarily target domain-specific extraction, confidence validation, or knowledge system extension rather than systematic standardization of extracted relationships to established ontological frameworks.
+chandak2023building emphasized the importance of standardized schemas and quality control in its PrimeKG framework by integrating heterogeneous data sources.
+RELATE applies these principles to the RE layer, enabling automated standardization of literature-derived relations for seamless integration into KGs such as PrimeKG, with ensured schema compliance and semantic interoperability.
+Recent work in LLM-based RE has explored various validation approaches.
+RelCheck (ourekouch2025relcheck) addresses low-confidence predictions from pretrained models using ontology-guided LLM validation, though it focuses on confidence rather than standardization.
+SPIREX (cavalleri2024spirex) applies schema-constrained prompts with graph machine learning validation but restricts itself to RNA-specific domains.
+Graph-augmented approaches include RAG-enhanced LLMs for automated ontology extension (georgakopoulos2025text), structure-oriented RAG for KG-LLM co-learning (yang2024knowledge), integrated GNNs with LLM-generated support documents (dong2024graph), and LLM-driven ontology enrichment pipelines (kollapally2025ontology).
+However, these approaches primarily target domain-specific extraction or confidence validation rather than systematic standardization of extracted relationships to established ontological frameworks.
 Multi-channel neural architectures combine textual and graph structures but focus primarily on structural features rather than systematic predicate mapping.
-Caufield et al.
-(2024) applied structured prompt interrogation for ontology-grounded annotations from scientific publications, yet their focus remained on entity linking rather than comprehensive relation standardization.
-Even enhancements to systems such as SemRep (Ming et al., 2024) that leverage contrastive learning operate within existing predicate frameworks rather than enabling mapping to broader ontological schemas.
+caufield2024structured used structured prompts for ontology-grounded annotations, focusing on entity linking rather than full relation standardization, while enhancements to SemRep (ming2024enhancing) remain within existing predicate frameworks.
+While cao2024automatic built an end-to-end system for extracting knowledge about rare diseases, their work addresses only 6 relation types in a narrow domain.
+More fundamentally, no existing framework systematically bridges free-text biomedical RE with standardized ontological predicates.
+RELATE addresses this critical gap by mapping free-text relations to standardized ontology predicates, handling both narrow (ChemProt’s 9 predicates) and broad (Biolink’s 284 predicates) ontologies, ensuring semantic interoperability and enabling seamless integration into biomedical KGs.
 3 RELATE Methodology
@@
 The first stage of our pipeline (Alg.
-1) extracts predicates with positive and negative descriptors to encode domain knowledge, and transforms them into a searchable embedding database suitable for relation mapping.
-We begin by collecting all predicate definitions from biomedical ontologies such as the Biolink Model toolkit (Unni et al., 2022) or ChemProt (Krallinger et al., 2017), thereby constructing a base vocabulary of biomedical relationships.
+1) extracts ontology predicates with positive and negative descriptors to encode domain knowledge, and transforms them into a searchable embedding database suitable for relation mapping.
+We begin by collecting all predicate definitions from biomedical ontologies such as the Biolink Model toolkit (unni2022biolink) or ChemProt (DBLP:journals/biodb/LiSJSWLDMWL16), thereby constructing a base vocabulary of biomedical relationships.
 We will denote the set of all ontology predicates as $\mathcal{R}$.
@@
 The descriptor dataset $\mathcal{D}_{r}^{+}$ is constructed by gathering existing textual descriptions for each predicate $r$.
-For example, one descriptor for the predicate affects may be “describes an entity that has an effect on the state or quality of another existing entity”.
-These descriptions are sourced from multiple authoritative biomedical sources.
+For example, one descriptor for the predicate affects may be “describes an entity that has an effect on the state or quality of another existing entity.” These descriptions are sourced from multiple authoritative biomedical sources.
 In the Biolink example, this includes official documentation and standardized vocabularies, while in the ChemProt case, online relation category descriptions suffice.
-The distributions of the predicate descriptors are presented in Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal.
+The distributions of the predicate descriptors are presented in Tables LABEL:tab:top_predicates_distribution_chemprot and LABEL:tab:top_predicates_distribution_heal.
 These aggregated descriptions form the positive dataset $\mathcal{D}_{r}^{+}$ and capture the semantic diversity with which biomedical relationships appear in scientific literature.
 To complement the positive dataset, we introduce a novel strategy for generating negative descriptors.
-For each descriptor in $D_{r}^{+}$, an LLM produces natural, semantically coherent negations using a structured prompt (Figure LABEL:fig:negation_prompt).
+For each descriptor in $D_{r}^{+}$, an LLM produces a natural, semantically coherent negation using a structured prompt (Figure LABEL:fig:negation_prompt).
 For example, the earlier descriptor becomes “describes an entity that does not have an effect on the state or quality of another existing entity.” These negated descriptors are stored in a separate dictionary $D_{r}^{-}$ to serve as contrastive examples that help the embedding model distinguish valid from invalid predicate mappings.
@@
 affects $\to$ affects_NEG) to maintain clear separation from positive instances.
-Finally, every descriptor in $D^{+}_{r}$ and $D^{-}_{r}$ is transformed into $d$-dimensional vector representations using the embedding model $f_{\theta}$.
+Finally, every descriptor in $D^{+}_{r}$ and $D^{-}_{r}$ is transformed into a $d$-dimensional vector representation using the embedding model $f_{\theta}$.
 Instead of aggregating these descriptor embeddings to a single predicate embedding, we store all descriptor embeddings, $\mathcal{V}_{r}$, in a searchable database for the next stage.
-To address limitations of general-purpose embedding models (e.g., nomic-text-embed or bge-m3) in biomedical domains, we introduce an (optional) hybrid retrieval enhancement based on SapBERT (Liu et al., 2021), a model specialized in biomedical entity linking.
+To address limitations of general-purpose embedding models (e.g., nomic-text-embed or bge-m3) in biomedical domains, we introduce an (optional) hybrid retrieval enhancement based on SapBERT (liu2021self), a model specialized in biomedical entity linking.
 In this setup (Figure LABEL:fig:workflowplus), SapBERT was fine-tuned on $D^{+}_{r}$ and $D^{-}_{r}$ to generate a second set of embeddings $\mathcal{V}^{\prime}_{r}$.
-While SapBERT has demonstrated effectiveness in biomedical entity alignment (Liu et al., 2021), this work presents its first application to relation extraction through predicate embedding enhancement.
+While SapBERT has demonstrated effectiveness in biomedical entity alignment, this work presents its first application to RE through descriptor embedding enhancement.
 Our hybrid approach leverages SapBERT’s biomedical domain knowledge to improve relation semantic similarity beyond general-purpose embedding models.
@@
 {{"negation_of_the_descriptor_text": "negated version" or "NOT ENOUGH INFORMATION"}}
-Figure 2: Negation generation prompt used in Stage 3.1 for creating negative descriptors from positive descriptors.
+Figure 2: Negation generation prompt used in Stage 1 (Section 3.1) for creating negative descriptors from positive descriptors.
 3.2 Stage 2: Similarity-Based Retrieval
 For every predicate $r\in\mathcal{R}$, Stage 1 has produced positive and negative descriptor dictionaries $\mathcal{D}_{r}^{+}$, $\mathcal{D}_{r}^{-}$ and corresponding descriptor embedding vectors $\mathcal{V}_{r}$.
-We now leverage these data to retrieve candidate ontology predicates for incoming extracted relations.
+We now leverage these data to retrieve candidate ontology predicates for incoming relations extracted by an upstream RE system, which could be any general-purpose LLM.
 The extracted relations come in a quadruple $(s,o,T,a)$: $s$ subject entity, $o$ object entity, $T$ the free-form relation text, and $a$ the abstract context from which the relation was identified.
-For example, $T$ could be “increases level of,” which does not exist in Biolink ontology in this format.
-First, free-form relation text $T$ is embedded to query vector $q=f_{\theta}(T)$.
+For example, in Figure LABEL:fig:workflow, $T$ is “decreases transporting activities of,” which does not exist in Biolink ontology $\mathcal{R}$ in this format.
+First, free-form relation text $T$ is embedded to a query vector $q=f_{\theta}(T)$.
 We use the same model $f_{\theta}$ from Stage 1, ensuring semantic consistency between query relations and stored ontology predicates.
@@
 By narrowing the search space from hundreds of ontology predicates to a small candidate set, this stage provides an efficient yet semantically informed mechanism for predicate selection.
-At this point, however, the search operates without the benefit of contextual information from the original biomedical abstract, focusing only on lexical and semantic similarity.
+At this point, however, the search operates without the benefit of contextual information from the original biomedical abstract $a$, focusing only on lexical and semantic similarity.
 3.3 Stage 3: Contextual Refinement
@@
 - If the original relationship implies negation (e.g., "does not cause"), select the matching base predicate, but set "negated" to "True".
-- If no match exists, return ‘"mapped_predicate": "none"‘.
+- If no match exists, return ‘"mapped_predicate": "none"’.
 Respond with ONLY this JSON object:
-{{"mapped_predicate": "one of the predicate keys or ’none’", "negated": "True" or "False"}}
-Figure 3: Contextual reranking prompt used in Stage 3 for LLM-based predicate selection with explicit negation handling.
+{{"mapped_predicate": "one of the predicate keys or ‘none’", "negated": "True" or "False"}}
+Figure 3: Contextual reranking prompt used in Stage 3 (Section 3.3) for LLM-based predicate selection with explicit negation handling.
 The final stage addresses the key limitation of similarity-based retrieval: its reliance on context-free embeddings.
 While Stage 2 efficiently narrows the predicate space to a small candidate set, it does not incorporate the biomedical context in which the relation was expressed.
-Stage 3 introduces contextual refinement, where an LLM evaluates and re-ranks the top-$k$ candidates against the full relation context.
+Stage 3 introduces contextual refinement, where an LLM evaluates and reranks the top-$k$ candidates against the full relation context.
 The LLM is given structured prompt (Figure LABEL:fig:reranking_prompt) along with the following information: the top candidate predicates retrieved by similarity-based search and the full RE quadruple $(s,o,T,a)$ including the abstract context.
 By explicitly grounding the candidate predicates in this richer context, the LLM can better assess which ontology predicate is most appropriate for the specific biomedical scenario.
-The LLM then performs re-ranking of the candidate set, selecting the single predicate that best aligns with the domain context.
+The LLM then performs reranking of the candidate set, selecting the single predicate that best aligns with the domain context.
 Importantly, the system includes a NONE option, allowing the model to reject all candidates if none provide a semantically valid match.
@@
 The output of Stage 3 is a contextually validated predicate $r^{*}$ that integrates both semantic similarity and biomedical context.
-This refinement ensures that only relationships consistent with both the ontology and the source text are admitted into the knowledge graph, thereby improving accuracy and reducing false positives.
+This refinement ensures that only relationships consistent with both the ontology and the source text are admitted into the KG, thereby improving accuracy and reducing false positives.
 4 Experiments
-[…]
-tab:chemprot_overall LLM Embedding Exact Match a@1 a@3 a@5 a@10 MRR MedGemma 27B nomic 0.464 0.528 0.686 0.744 0.755 0.612 MedGemma 27B bge-m3 0.467 0.468 0.686 0.788 0.812 0.590 MedGemma 27B nomic + SapBERT 0.520 0.100 0.328 0.688 0.914 0.300 MedGemma 27B bge + SapBERT 0.516 0.100 0.321 0.691 0.940 0.302
-Table 1: RELATE performance on ChemProt. Exact Match: accuracy of RELATE; a@$k$: accuracy in top-$k$ candidates before LLM-based refinement in Stage 3 of RELATE; MRR: Mean reciprocal rank.
-We assess RELATE on two independent datasets with different ontology models, as well as four different embedding model configurations.
+We assess RELATE on two independent datasets with different ontology models, as well as four different embedding model configurations. See Appendix D for example descriptors in both datasets.
 4.1 Datasets
-The ChemProt dataset (Krallinger et al., 2017) provides annotated chemical-protein relations representing diverse pharmacological and biochemical relationship types, including regulators, modulators, agonists, antagonists, and substrate interactions.
+The ChemProt dataset (DBLP:journals/biodb/LiSJSWLDMWL16) provides annotated chemical-protein relations representing diverse pharmacological and biochemical relationship types, including regulators, modulators, agonists, antagonists, and substrate interactions.
 We treat the information in ChemProt as ground truth abstract-predicate pairs for evaluation purposes.
 Standard metrics in information retrieval are used, such as exact match accuracy, accuracy@$k$, and Mean Reciprocal Rank (MRR).
-Recall from Figure LABEL:fig:reranking_prompt that Stage 3 of RELATE only outputs its top choice predicate $r^{*}$ or NONE, and does not return a fully re-ranked or re-ordered list of $k$ candidates.
-Therefore, exact match accuracy is calculated with respect to the final output of RELATE, $r^{*}$, but accuracy@$k$ and MRR are calculated using the top-$k$ candidates after Stage 2, before LLM-based Stage 3.
-This will allow us to analyze the effect of LLM-based refinement in isolation as well.
+Recall from Figure LABEL:fig:reranking_prompt that Stage 3 of RELATE only outputs its top choice predicate $r^{*}$ or NONE, and does not return a fully reranked or reordered list of $k$ candidates.
+Therefore, exact match accuracy is calculated with respect to the final output of RELATE ($r^{*}$), but accuracy@$k$ and MRR are calculated using the top-$k$ candidates after Stage 2, before LLM-based Stage 3.
+This allows us to analyze the impact of LLM-based refinement in isolation.
 Next, we apply RELATE to 2,400 PubMed abstracts from the HEAL Project [^4] https://heal.nih.gov targeting opioid use disorder research.
-This real-world dataset reflects the practical application scenario where relationships extracted from literature must be mapped to standardized Biolink predicates for integration into a large biomedical KG [^5] Name hidden during double blind review process..
+This real-world dataset reflects the practical scenario in which extracted relationships must be mapped to standardized Biolink predicates for integration into a large biomedical KG [^5] https://robokop.renci.org/.
 Unlike ChemProt, the HEAL abstracts lack ground-truth annotations, requiring qualitative assessment of RELATE’s outputs.
 Importantly, this corpus captures the complexity and noisiness of real biomedical literature.
-In contrast to benchmark datasets, these abstracts often contain methodological statements or negative scientific findings (e.g., drug does not affect condition, gene not associated with disease).
-Such cases underscore the necessity of RELATE’s rejection and negation-handling capabilities to ensure only semantically appropriate biomedical relationships enter the KG.
-As part of the integration plan, the abstract authors and Subject Matter Experts (SMEs) are being recruited to systematically evaluate the quality and reliability of the extracted triples.
+In contrast to benchmark datasets, these abstracts often contain methodological statements or negative scientific findings (e.g., “drug does not affect condition,” “gene not associated with disease”).
+Such cases underscore the necessity of rejection and negation-handling to ensure only semantically appropriate biomedical relationships enter the KG.
 4.2 LLM and Embedding Models
@@
 The last two demonstrated greater consistency and fewer hallucinations compared to the other models.
-Since those two models had very similar performances, we only report results using MedGemma in this work.
+Since those two models had very similar performances, we only report results using MedGemma (27B) in this work.
 However, note that any other LLM could be substituted to reproduce the pipeline.
-MedGemma-27B is Google’s medical language model (Sellergren et al., 2025), trained exclusively on medical text and optimized for efficient inference through architectural improvements in performance and computational efficiency.
+MedGemma (27B) is Google’s medical language model (sellergren2025medgemma), trained exclusively on medical text and optimized for efficient inference through architectural improvements in performance and computational efficiency.
+To evaluate RELATE’s pipeline contribution beyond raw LLM capability, we implemented a direct classification baseline.
+This baseline uses MedGemma (27B) to classify ChemProt relations by providing the same input quadruple (subject, object, relation text, abstract) along with the list of 9 ChemProt predicates.
+This represents a common approach in LLM-based relation extraction and provides a fair comparison since it uses the same LLM and has access to the same predicate definitions.
 As with the LLMs, we also experimented with several embedding approaches.
 While we prioritize accessibility through ollama-based open-access models, we tested OpenAI’s text-embed-large models and others with 8,192-token context windows that provide general-purpose semantic understanding.
-These include nomic-embed-text (Zach et al., ), a 768-dimensional embedding model, and bge-m3, a 1,024-dimensional multilingual embedding model for diverse textual contexts (Chen et al., 2024b).
+These include nomic-embed-text (nussbaumnomic), a 768-dimensional embedding model, and bge-m3, a 1,024-dimensional multilingual embedding model for diverse textual contexts (chen2024bge).
 Since nomic-embed-text and bge-m3 achieved comparable performance, we report only the results obtained with the 768-dimensional nomic-embed-text in this work.
-For (optional) hybrid retrieval, we experimented with adding fine-tuned SapBERT (Liu et al., 2021).
+For (optional) hybrid retrieval, we experimented with adding fine-tuned SapBERT (liu2021self).
 We initialized the model with BiomedNLP-PubMedBERT-base-uncased-abstract -fulltext and optimized it using a contrastive learning framework.
@@
 We used mixed-precision training to reduce computational overhead and set a fixed random seed for reproducibility.
+[…]
+tab:chemprot_overall Embedding Model Stage 2 (Candidate Retrieval) Post-Refinement a@1 a@3 a@5 a@10 MRR Exact Match (a@1) nomic 0.528 0.686 0.744 0.755 0.612 0.464 bge-m3 0.468 0.686 0.788 0.812 0.590 0.467 nomic + SapBERT 0.100 0.328 0.688 0.914 0.300 0.520 bge-m3 + SapBERT 0.100 0.321 0.691 0.940 0.302 0.516 Baseline (no embedding) - - - - - 0.220
+Table 1: RELATE performance on ChemProt.
+Stage 2 (Candidate Retrieval): a@$k$ = accuracy in top-$k$ candidates before LLM-based refinement; MRR = mean reciprocal rank.
+Post-Refinement: Exact Match = final accuracy after LLM-based refinement; Baseline = direct LLM classification without RELATE’s retrieval-refinement pipeline.
+LLM used is MedGemma (27B) for all.
 5 Results and Discussions
+RELATE Pipeline is Necessary.
+To validate RELATE’s pipeline design, we compared against direct LLM classification where MedGemma (27B) selects from the 9 ChemProt predicates without retrieval-refinement.
+The baseline achieves only 22.0% exact match (Table LABEL:tab:chemprot_overall), compared to RELATE’s 52.0% (136% improvement).
+This demonstrates that structured retrieval outperforms asking an LLM to select directly from a predicate list.
+The baseline’s poor performance reflects the difficulty of simultaneous predicate understanding and selection without semantic guidance from embedding-based retrieval.
+The exact match accuracy of 52% reflects the difficulty of mapping to fine-grained ontology predicates.
+However, this metric may understate practical utility.
+Manual analysis of the misclassifications show semantically acceptable near-misses, such as mapping to agonist instead of modulator.
+When allowing such semantically close alternatives, accuracy increases to approximately 70-80%.
 LLM-based Refinement is Necessary.
@@
 These results highlight a trade-off: baseline embeddings rank the correct answer more accurately at the top, while SapBERT embeddings broaden coverage and guarantee inclusion within candidate sets.
-RELATE resolves this through its LLM-based refinement stage, which reranks SapBERT’s expanded candidate lists in context.
-This allows RELATE to recover early precision while preserving high recall, transforming broad retrieval into precise ontology-constrained mappings.
-In short, SapBERT enhances retrieval coverage, but RELATE’s LLM-based refinement is essential to convert this coverage into accurate, ontology-aligned relations.
+While selecting the top-ranked candidate from embeddings without LLM refinement would yield higher exact match on this benchmark, such an approach assumes embedding similarity reliably ranks correct predicates first across all datasets and relation types.
+This assumption cannot be guaranteed a priori.
+RELATE’s design prioritizes robustness through SapBERT’s comprehensive retrieval combined with LLM-based contextual reranking, which serves as a buffer against cases where embedding similarity alone performs poorly.
+This allows RELATE to recover early precision while preserving high recall, transforming broad retrieval into precise ontology-constrained mappings regardless of how initial embeddings behave on new corpora.
 When Does RELATE Reject Candidates?
@@
 Out of 2,400 PubMed abstracts from the HEAL project, 10 relationships were appropriately rejected as unsuitable for Biolink predicate mapping (rejection rate:  0.4%).
-Analysis of these rejections reveals contextual understanding that distinguishes between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
+Manual analysis of these rejections reveals contextual understanding that distinguishes between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
 The rejected relationships fell into distinct categories reflecting non-biomedical associations: genetic nomenclature specifications (3 cases), clinical procedures and documentation (4 cases, including “Clinicians prescribe Opioid” and “Women on BUP had documented Opioid misuse”), methodological relationships (2 cases, such as “Anisomycin used to detect GNPTG”), and specialized reporting contexts (1 case).
@@
 These rejections occurred despite high embedding similarity scores, demonstrating the reranker’s comprehensive contextual analysis.
-The low rejection rate (0.4%) indicates that the majority of extracted relationships represent valid biomedical associations while maintaining strict quality standards for edge cases requiring specialized handling.
-Chemprot analysis also reveals potential false rejections where high similarity scores failed to produce mappings.
-Cases like “loperamide modifies corticotropin-releasing hormone actio“ (0.661 score for “modulator“) and “ammonium sulfate reduces CBG concentrations“ (0.632 for “downregulator“) were rejected despite appropriate candidates.
+The low rejection rate (0.4%) suggests that the majority of extracted relationships represent valid biomedical associations while maintaining strict quality standards for edge cases requiring specialized handling.
+All 10 rejections were deemed appropriate (100% precision).
+On ChemProt, there were 5 correct rejections out of 18 total rejections.
+These false rejections occur where, despite high similarity candidate scores, RELATE failed to produce mappings.
+Cases like “loperamide modifies corticotropin-releasing hormone action” (0.661 score for modulator) and “ammonium sulfate reduces CBG concentrations” (0.632 for downregulator) were rejected despite appropriate candidates.
 While some discrepancies reflect stricter semantic criteria, others suggest overly conservative contextual reranking, highlighting evaluation challenges with traditional benchmarks.
+These contrasting patterns reflect differences in semantic alignment between upstream extraction and ontology predicates across datasets.
+See more details in Appendix C.2.
 When Does RELATE Choose Negated Candidates?
@@
 The remaining quarter of cases presented genuine linguistic challenges.
-Consider phrases like “remained impaired despite treatment” or “not necessary for acute insulin-induced uptake”.
-These require contextual understanding that goes well beyond pattern matching.
+Consider phrases like “remained impaired despite treatment” or “not necessary for acute insulin-induced uptake.” These require contextual understanding that goes well beyond pattern matching.
 The system had to recognize that “fails to affect” and “unable to deliver” carry negative semantic weight, even when the word “not” never appears.
+See more details in Appendix C.3.
 Perhaps most importantly, RELATE maintained conceptual accuracy throughout this process.
@@
 Clinical terminology refinement was most prominent, with 28 cases shifting from associated with: increased likelihood of to predisposes to condition.
-Domain specificity enhancement appeared in 12 cases where generic associated with became gene associated with condition for relationships like “DRD4 associated with Heroin Addiction,” recognizing genetic contexts require more informative predicates.
+Domain specificity enhancement appeared in 12 cases where generic associated with became gene associated with condition for relationships like “DRD4 associated with Heroin Addiction,” recognizing that genetic contexts require more informative predicates.
 Evidence-based caution moderated 12 strong causal claims, changing causes to contributes to for relationships like “MDMA causes Depression,” reflecting the complex multifactorial causation typical in biomedicine, where direct causality claims often overstate evidence strength.
@@
 These systematic patterns demonstrate that SapBERT’s biomedical entity embeddings guide predicate selection toward clinically appropriate, evidence-aware terminology that better reflects expert knowledge of biomedical relationship complexity.
+Limitations.
+RELATE’s performance depends on the quality of upstream entity and relation extraction.
+While we do not assume perfect extraction, the input quadruples must be reasonably accurate.
+Errors in entity boundaries, entity type misclassifications, or poorly extracted relation phrases will propagate through the standardization pipeline.
+In production KG construction, additional methods address these entity errors.
+RELATE’s modular design allows it to benefit from improvements in upstream extraction systems without requiring system redesign.
 6 Conclusion
 We introduced RELATE, a three-stage pipeline for automated ontology-constrained predicate mapping in biomedical relation extraction.
-By combining ontology-driven preprocessing, efficient vector search, and context-aware LLM reranking, RELATE systematically converts free-text relations into ontology-aligned, interoperable knowledge graph edges.
-Experiments on ChemProt demonstrate the trade-offs between early precision and retrieval depth, while application to 2,400 HEAL Project abstracts illustrates RELATE’s ability to handle noisy real-world biomedical literature, including identification of negated assertions and appropriate rejection of non-biomedical associations.
-Together, these results establish a foundation for scalable, ontology-constrained relation extraction and highlight a path forward for enriching biomedical knowledge graphs with high-quality, standardized information.
-Future work includes improving computational overhead, designing a scalable evaluation that combines expert judgment with automated consistency checks, and integration into the existing large biomedical KG (ROBOKOP).
+By combining ontology-driven preprocessing, efficient vector search, and context-aware LLM-based reranking, RELATE systematically converts free-text relations into ontology-aligned and interoperable KG edges.
+Experiments on ChemProt demonstrate the trade-offs between early precision and retrieval depth, while application to 2,400 HEAL Project abstracts illustrates RELATE’s ability to handle noisy real-world biomedical literature, including detecting negated assertions and appropriately rejecting non-biomedical associations.
+Together, these results establish a foundation for scalable, ontology-constrained relation extraction and highlight a path forward for enriching biomedical KGs with high-quality, standardized information.
+Future work includes improving computational overhead, incorporating further top-$k$ reordering and evaluation in the refinement stage, designing a scalable evaluation that combines expert judgment with automated consistency checks, and integration into the existing large biomedical KG (ROBOKOP).
+As part of the integration plan, the HEAL abstract authors and domain experts are being recruited to systematically evaluate the quality and reliability of RELATE and the LLM-extracted quadruples that serve as its inputs.
 […]
-We thank the LitCoin project partners for their immense contributions, the HEAL initiative consortium for providing access to opioid research abstracts.
+We thank the LitCoin project partners for their immense contributions, and the HEAL initiative consortium for providing access to opioid-related research abstracts.
 Also, we acknowledge the RENCI’s Data Analytics group and Knowledge Graph Working Group for constructive reviews.
-This research was supported by NIH #75N95023C00032 (LITCOIN)
-Appendix A Relation Extraction Pipeline Stages Algorithms
-Input: Biolink Model YAML $\mathcal{B}$, Biolink Toolkit $\mathcal{T}$, LLM $\mathcal{M}$, Embedding model $f_{\theta}$
-Output: Embedding store $\mathcal{V}$, Positive descriptors $\mathcal{D}^{+}$, Negative descriptors $\mathcal{D}^{-}$
+This research was supported by NIH #75N95023C00032 (LITCOIN).
+Appendix A Clinical and Research Applications
+A.1 RELATE’s Role and Significance
+RELATE is a relation standardization component that ensures relations extracted from biomedical text are accurately mapped to standardized ontology predicates.
+This standardization is foundational for downstream knowledge graph applications.
+In literature-based discovery, standardized predicates enable systematic querying across diverse sources.
+A researcher investigating drug repurposing can query for all compounds that upregulate proteins downregulated in a specific disease, confident that the upregulator predicate maintains consistent semantics across thousands of papers.
+Without standardization, the same relationship might appear as “increases expression,” “upregulates,” or “enhances activity,” requiring separate queries and potentially missing relevant findings.
+Biomedical knowledge graph systems such as ROBOKOP require ontology-aligned relations for reliable reasoning and inference.
+RELATE ensures that literature-derived evidence integrates seamlessly with structured knowledge bases, enabling complex queries that traverse multiple relationship types and data sources.
+Our evaluation on HEAL demonstrates how relations from distinct domains can be mapped to common Biolink predicates, while ChemProt enables unified knowledge graphs that integrate chemical interactions.
+A.2 Comparison with End-to-End Systems
+Unlike end-to-end systems such as AutoRD, which perform extraction, standardization, and graph construction, RELATE focuses specifically on standardization.
+This provides modularity, allowing RELATE to work with any upstream extraction system, and scalability, handling 284 Biolink predicates compared to 6 relation types in AutoRD.
+A.3 Broader Applicability
+While evaluated on biomedical corpora, RELATE generalizes to any domains with structured predicate vocabularies for computing embeddings.
+Appendix B RELATE Stages Algorithms
+Algorithms 1, 2, and 3 summarize the three stages of RELATE. When the optional SAPBERT workflow enhancement is on, RELATE follows the modified pipeline described in Figure LABEL:fig:workflowplus.
+Input: Biolink Model YAML $\mathcal{B}$, Biolink Toolkit $\mathcal{T}$, Embedding model $f_{\theta}$, LLM
 […]
 $\mathcal{D}^{+}\leftarrow\emptyset$;
-for each predicate mapping $r\in\mathcal{R}$ do
- $\text{predicate}\leftarrow r.\text{predicate}$;
- $\text{descriptions}\leftarrow$ GatherDescriptions(predicate)($\text${predicate});
- $\mathcal{D}^{+}[\text{predicate}]\leftarrow\text{descriptions}$;
+for each predicate $r\in\mathcal{R}$ do
+ $\text{descriptions}\leftarrow$ GatherDescriptions(r)(r);
+ $\mathcal{D}^{+}[r]\leftarrow\text{descriptions}$;
 end for
@@
 for each predicate $p\in\mathcal{D}^{+}$ do
- $\text{pos\_descriptions}\leftarrow\mathcal{D}^{+}[p]$;
  $\text{neg\_descriptions}\leftarrow\emptyset$;
- for each description $d\in\text{pos\_descriptions}$ do
+ for each description $d\in\mathcal{D}^{+}[p]$ do
     $\text{prompt}\leftarrow$ CreateNegationPrompt(d)(d);
-    $d^{neg}\leftarrow\mathcal{M}(\text{prompt})$;
-    $\text{neg\_descriptions}\leftarrow\text{neg\_descriptions}\cup\{d^{neg}\}$;
+    $d^{-}\leftarrow LLM(\text{prompt})$;
+    $\text{neg\_descriptions}\leftarrow\text{neg\_descriptions}\cup\{d^{-}\}$;
    end for
- $\text{neg\_key}\leftarrow p+\text{"\_NEG"}$;
+ $\text{neg\_key}\leftarrow p+\text{``\_NEG"}$;
  $\mathcal{D}^{-}[\text{neg\_key}]\leftarrow\text{neg\_descriptions}$;
@@
  $\text{combined\_text}\leftarrow$ CombineDescriptions(descriptions)($\text${descriptions});
- $\mathbf{v}_{p}\leftarrow f_{\theta}(\text{combined\_text})$;
- $\mathcal{V}\leftarrow\mathcal{V}\cup\{(\text{label}:p,\text{embedding}:\mathbf{v}_{p})\}$;
+ $v\leftarrow f_{\theta}(\text{combined\_text})$;
+ $\mathcal{V}\leftarrow\mathcal{V}\cup\{(\text{label}:p,\text{embedding}:v)\}$;
 end for
-return $\mathcal{V},\mathcal{D}^{+},\mathcal{D}^{-}$;
+return Embedding store $\mathcal{V}$, Positive descriptors $\mathcal{D}^{+}$, Negative descriptors $\mathcal{D}^{-}$;
 Algorithm 1 Ontology Preprocessing
-Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Predicate embeddings $\mathcal{V}$, Embedding model $f_{\theta}$, Top-k parameter $k$
-Output: $\mathcal{C}$
+Input: Loose quadruples $\{(s_{i},o_{i},T_{i},a_{i})\}$, Predicate embeddings $\mathcal{V}$, Embedding model $f_{\theta}$, Parameter $k$
 // Relationship Embedding Phase
-$\mathcal{R}_{embed}\leftarrow\emptyset$;
-for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
- $\mathbf{v}_{r_{i}}\leftarrow f_{\theta}(r_{i})$;
- $\mathcal{R}_{embed}\leftarrow\mathcal{R}_{embed}\cup\{(i,\mathbf{v}_{r_{i}})\}$;
+$\mathcal{Q}\leftarrow\emptyset$;
+for each quadruples $(s_{i},o_{i},T_{i},a_{i})$ do
+ $q_{i}\leftarrow f_{\theta}(T_{i})$;
+ $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(i,q_{i})\}$;
 end for
@@
 $\mathcal{C}\leftarrow\emptyset$;
-for each $(i,\mathbf{v}_{r_{i}})\in\mathcal{R}_{embed}$ do
+for each $(i,q_{i})\in\mathcal{Q}$ do
  $\text{similarities}\leftarrow\emptyset$;
- for each $(\text{label}_{j},\mathbf{v}_{j})\in\mathcal{V}$ do
-    $\text{sim}_{j}\leftarrow\cos(\mathbf{v}_{r_{i}},\mathbf{v}_{j})=\frac{\mathbf{v}_{r_{i}}\cdot\mathbf{v}_{j}}{\|\mathbf{v}_{r_{i}}\|\|\mathbf{v}_{j}\|}$;
-    $\text{similarities}\leftarrow\text{similarities}\cup\{(\text{label}_{j},\text{sim}_{j})\}$;
+ for each $(p,v)\in\mathcal{V}$ do
+    $\text{sim}\leftarrow\cos(q_{i},v)$;
+    $\text{similarities}\leftarrow\text{similarities}\cup\{(p,\text{sim})\}$;
    end for
- $\text{sorted}\leftarrow$ SortDescending(similarities)($\text${similarities});
- $\text{candidates}_{i}\leftarrow\text{sorted}[1:k]$;
- $\mathcal{C}\leftarrow\mathcal{C}\cup\{(i,\text{candidates}_{i})\}$;
+ $\mathcal{C}[i]\leftarrow$ Top-kk(similarities)($\text${similarities});
 end for
-return $\mathcal{C}$;
+return Candidate sets $\mathcal{C}$;
 Algorithm 2 Similarity-Based Retrieval
-Input: Loose triples $\mathcal{T}=\{(s_{i},o_{i},r_{i},a_{i})\}_{i=1}^{n}$, Candidate sets $\mathcal{C}$, LLM $\mathcal{M}$
-Output: Mapped triples $\mathcal{T}^{*}=\{(s_{i},o_{i},r_{i}^{*},a_{i})\}_{i=1}^{n}$
-$\mathcal{T}^{*}\leftarrow\emptyset$;
-for each triple $(s_{i},o_{i},r_{i},a_{i})\in\mathcal{T}$ do
- $\text{candidates}_{i}\leftarrow\mathcal{C}[i]$;
+Input: Loose quadruples $\{(s_{i},o_{i},T_{i},a_{i})\}$, Candidate sets $\mathcal{C}$, LLM
+$\mathcal{A}\leftarrow\emptyset$;
+for each quadruple $(s_{i},o_{i},T_{i},a_{i})$ do
 […]
- // Prompt Construction
- $\text{prompt}\leftarrow$ ”Given the biomedical relationship context:”;
- $\text{prompt}\leftarrow\text{prompt}+\text{" Subject: "}+s_{i}+\text{" Object: "}+o_{i}$;
- $\text{prompt}\leftarrow\text{prompt}+\text{" Relationship: "}+r_{i}+\text{" Abstract: "}+a_{i}$;
- $\text{prompt}\leftarrow\text{prompt}+\text{" Select best Biolink predicate or NONE:"}$;
- for $j=1$ to $|\text{predicate\_list}|$ do
-    $\text{prompt}\leftarrow\text{prompt}+j+\text{". "}+\text{predicate\_list}[j]$;
-   end for
- $\text{prompt}\leftarrow\text{prompt}+(|\text{predicate\_list}|+1)+\text{". NONE"}$;
+[…]
  // LLM Selection and Validation
- $\text{response}\leftarrow\mathcal{M}(\text{prompt})$;
-[…]
- if $r_{i}^{*}=\text{"NONE"}$ or $r_{i}^{*}\notin\text{predicate\_list}$ then
+ $r_{i}^{*}\leftarrow LLM(\text{prompt})$;
+ if $r_{i}^{*}=\texttt{NONE}$ or $r_{i}^{*}\notin\text{predicate\_list}$ then
     $r_{i}^{*}\leftarrow\emptyset$;
    end if
- $\mathcal{T}^{*}\leftarrow\mathcal{T}^{*}\cup\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
+ $\mathcal{A}\leftarrow\mathcal{A}\cup\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
 end for
-return $\mathcal{T}^{*}$;
+return Mapped quadruples $\mathcal{A}=\{(s_{i},o_{i},r_{i}^{*},a_{i})\}$;
 Algorithm 3 Contextual Refinement
-Appendix B SAPBERT Workflow Enhancement
 […]
 fig:workflowplus Refer to caption
-Figure 4: SapBERT Augmented System workflow diagram showing the three-stage pipeline as in Figure LABEL:fig:workflow: (1) Ontology Preprocessing, (2) Similarity-Based Retrieval, and (3) Contextual Refinement.
-Just as in Figure LABEL:fig:workflow, stage 3.1 changes only if the ontology schema or the LLM and embedding Model changes.
-However, the one-stage preprocessing of raw ontology predicate data involves two embedding generations- LLM-generated embedding and SapBERT finetuned stored embedding.
-Also, given a raw input triple with abstract context, the frequently executed stage—similarity-based retrieval and contextual refinement—encompasses the input relation’s dual embedding transformation, dual similarity search on the pair of stored embeddings, merger of top-k candidates from each search, and final contextual reranking.
-Appendix C Predicate and Descriptor Distribution
+Figure 4: SapBERT-augmented RELATE pipeline as in Figure LABEL:fig:workflow.
+The ontology preprocessing now involves both LLM-generated embeddings and SapBERT finetuned embeddings.
+Additional changes include dual embedding of the relation text, dual similarity searches on the stored embeddings, and merging of the two top-$k$ candidates.
+Appendix C Additional Analysis
+C.1 ChemProt Pre/Post Refinement Accuracies: Error Attribution Analysis
+Our results revealed an unexpected pattern: Stage 2 (Section 3.2) retrieval achieved 52.8% accuracy@1, but after Stage 3 (Section 3.3) refinement, final accuracy dropped to 46.6% (Table LABEL:tab:chemprot_overall).
+To understand this decline, we analyzed the 6.2% set failures and attributed them to specific causes.
+C.1.1 Stage 3 Impact Analysis
+We identified all cases where Stage 3 (Section 3.3) changed Stage 2’s (Section 3.3) @1 prediction and categorized them by outcome:
+  1. 1.
+    Stage 3 harmful changes (146 cases): Stage 3 changed a correct @1 prediction to an incorrect answer
+  2. 2.
+    Stage 3 beneficial changes (103 cases): Stage 3 rescued failures by selecting the correct answer from ranks 2-10
+The net impact was -43 cases (146 harmful - 103 beneficial), accounting for the observed 6.2% accuracy decline.
+This analysis focuses on the 146 harmful changes to understand why Stage 3 overrode correct predictions.
+C.1.2 Systematic Over-Specification Pattern
+Detailed examination of the 146 harmful changes revealed highly concentrated error patterns rather than random failures (Table 2).
+A single transformation pattern in Stage 2 correctly predicted “downregulator” but Stage 3 changed it to “inhibits”; accounting for 50.7% of all harmful changes.
+The top five patterns captured 83.6% of errors, indicating systematic bias.
+Table 2: Stage 3 Over-Specification Error Patterns with Original Free-Text Relationship Variants
+Ground Truth Stage 3 Original Free-Text Variants Count %
+downregulator inhibits “inhibits” (68), “inhibits the activity of” (4), “inhibits secretion of” (1), “inhibits the activities of” (1) 74 50.7
+downregulator antagonist “blocks” (8), “inhibits” (6), “blockers of” (1), “is a poor inhibitor of” (1), “tyrosine kinase inhibitor” (1), “interferes with inhibitory action on” (1) 18 12.3
+downregulator inhibitor “is an inhibitor of” (3), “inhibits” (3), “inhibitor of” (2), “selectively inhibited” (1), “preferentially inhibits” (1), “inhibitory activity” (1), +4 more 15 10.3
+upregulator agonist “activates” (4), “induces expression of” (3), “stimulates” (2), “stimulates the transcription of” (1), “increases phosphorylation of” (1) 11 7.5
+upregulator activates “activates” (4) 4 2.7
+Top 5 patterns 122 83.6%
+upregulator modulator “activates” (1), “stimulates activity of” (1), “stimulates the activity of” (1) 3 2.1
+downregulator regulator “inhibits” (1), “is abolished by” (1), “reduces concentrations of” (1) 3 2.1
+downregulator blocks “blocks” (3) 3 2.1
+substrate regulator “impacts metabolism of” (2) 2 1.4
+Other patterns (6 additional) 12 8.2
+Total 146 100.0%
+Analysis of the original free-text relationships revealed the fundamental mechanism:
+  1. 1.
+    Stage 3 preserved mechanistic specificity from source text rather than maintaining ontology-level abstraction.
+In the dominant error pattern (downregulator → inhibits), 92% of cases (68/74) had “inhibits” as the original free-text relationship
+  2. 2.
+    Stage 2 (retrieval-based) correctly abstracted to the ontology-level category “downregulator,” matching ChemProt’s annotation guidelines.
+However, Stage 3 (LLM-based refinement) reintroduced the mechanistic specificity from the source text, selecting “inhibits” because that was the literal wording in the abstract.
+This pattern extended across error types: “blocks” appeared in 61% (11/18) of cases where Stage 3 incorrectly changed “downregulator” to “antagonist,” and “activates” appeared in 73% (8/11) of cases where Stage 3 changed “upregulator” to “agonist.”
+C.1.3 Interpretation
+These findings illuminate a fundamental tension between retrieval-based and generative approaches to relation extraction.
+The embedding-based retrieval (Stage 2 (Section 3.2)) successfully learned ontology-level abstractions from the training data, correctly mapping mechanistic phrases to categorical predicates.
+However, the LLM refinement (Stage 3 (Section 3.2)), despite being trained extensively on biomedical literature, demonstrated a preference for preserving the literal mechanistic terminology found in source text.
+The concentration of errors (84%) captured by just five patterns suggests that targeted interventions could recover most of the 6.2% accuracy loss:
+  1. 1.
+    Confidence-based bypass: Skip Stage 3 when Stage 2’s top-1 confidence exceeds a threshold (e.g., 0.8)
+  2. 2.
+    Ontology-constrained prompting: Explicitly instruct the LLM to maintain categorical abstraction levels
+  3. 3.
+    Restricted candidate sets: Only allow Stage 3 to choose from predicates at the same abstraction level as the top-k Stage 2 candidates
+C.2 Rejected Relations
+RELATE includes a rejection mechanism (NONE option, Section 3.3) to filter relations that cannot be reliably mapped to ontology predicates. Here, we analyze rejection patterns across both datasets.
+C.2.1 HEAL Dataset Rejections
+In the HEAL dataset of 2,400 abstracts, RELATE rejected 10 relationships (rejection rate: 0.4%).
+Manual review confirmed that all 10 rejections were appropriate (100% precision), filtering non-biomedical associations.
+The rejected relationships fell into distinct categories:
+  1. 1.
+    Genetic nomenclature specifications (3 cases): Relations describing genetic naming conventions rather than biological relationships (e.g., gene symbol designations)
+  2. 2.
+    Clinical procedures and documentation (4 cases): Administrative or procedural relationships such as “Clinicians prescribe Opioid” and “Women on BUP had documented Opioid misuse.”
+  3. 3.
+    Methodological relationships (2 cases): Experimental procedures such as “Anisomycin used to detect GNPTG.”
+  4. 4.
+    Specialized reporting contexts (1 case): Context-specific descriptions not representing biological relationships.
+These rejections demonstrate RELATE’s ability to distinguish between legitimate biomedical relationships and other types of associations commonly found in scientific literature.
+All rejections were semantically distant from Biolink predicates, making the NONE decision appropriate.
+C.2.2 ChemProt Dataset Rejections
+Analysis of the ChemProt test set (695 cases with ground truth labels) revealed 18 instances where RELATE selected NONE.
+Of these, only 5 were correct rejections, yielding a rejection precision of 27.8% (72% false rejection rate among rejected cases).
+The 13 incorrect rejections occurred despite appropriate candidate predicates being retrieved with moderate-to-high similarity scores during Stage 2 (Section 3.2). Examples include:
+  1. 1.
+    “loperamide modifies corticotropin-releasing hormone action” retrieved modulator with similarity score 0.661, yet was rejected in Stage 3.
+  2. 2.
+    “ammonium sulfate reduces CBG concentrations” retrieved downregulator with similarity score 0.632, yet was rejected in Stage 3.
+These patterns indicate that incorrect rejections stem from overly conservative contextual evaluation during LLM-based refinement in Stage 3 rather than retrieval failure in Stage 2.
+The system prioritizes precision by applying strict semantic criteria when assessing predicate-context fit, which reduces false positives but increases false rejections in ambiguous cases.
+C.3 Negation Handling
+We conducted a manual evaluation of negation detection on the HEAL dataset.
+Of 2,411 relationship extractions from 2,400 abstracts, we identified 123 (5.1%) negated relation texts through manual review.
+RELATE correctly flagged 74 as negated assertions, yielding:
+  1. 1.
+    Precision: 96.1% (74 true positives / 77 flagged as negated)
+  2. 2.
+    Recall: 60.2% (74 true positives / 123 actual negations)
+  3. 3.
+    F1-score: 74.2%
+The high precision indicates that when RELATE identifies a negation, it is almost always correct (only 3 false positives).
+The moderate recall indicates that approximately 40% of negated statements are missed (49 false negatives), typically involving:
+  1. 1.
+    Implicit negations (e.g., “failed to affect”, “unable to deliver”)
+  2. 2.
+    Comparative negations (e.g., “no significant difference”)
+  3. 3.
+    Complex syntactic constructions not captured by pattern matching
+The negation categories in the 123 manually identified cases were:
+  1. 1.
+    Explicit negations (“not”, “no”, “does not”): 90 cases (73%)
+  2. 2.
+    Implicit negations (“failed to”, “unable to”): 21 cases (17%)
+  3. 3.
+    Comparative negations (“no significant difference”): 12 cases (10%)
+RELATE’s negation handling maintains conceptual accuracy by identifying the underlying relationship type while flagging the negation polarity, enabling accurate representation of negative findings in KGs.
+Appendix D Predicate and Descriptor Distributions
+ChemProt defines 10 classes for chemical-protein interactions: 9 positive relationship types and 1 negative class (NOT/no interaction).
+We utilize all 9 positive relationship types, yielding 30 positive descriptors (Table LABEL:tab:top_predicates_distribution_chemprot).
+ChemProt’s NOT class, which indicates no interaction exists, is handled through RELATE’s explicit rejection (NONE) mechanism rather than treating NOT/no interaction as a mappable predicate.
+This design ensures that only valid biomedical relationships are standardized to ontology predicates while filtering non-relationships before KG integration.
+After the preprocessing stage (Section 3.1), the ontology database stores both positive predicates and descriptors and their LLM-negated versions.
+For ChemProt, this yields 18 stored predicates (9 positive + 9 NEG variants) with 60 total descriptor texts (30 positive + 30 negated).
+For Biolink, this yields 568 stored predicates (284 positive + 284 NEG variants) with 2,208 total descriptor texts (1,104 positive + 1,104 negated).
+Tables LABEL:tab:top_predicates_distribution_chemprot and LABEL:tab:top_predicates_distribution_heal show the descriptor distribution for the most frequent predicates.
+In ChemProt (Table LABEL:tab:top_predicates_distribution_chemprot), substrate and agonist together account for over one-third of all positive descriptors (36.67%).
+In Biolink (Table LABEL:tab:top_predicates_distribution_heal), the top 10 predicates account for 41.30% of all positive descriptors, with coexists with (16.12%) and related to (8.70%) being the most prevalent.
+The remaining 274 Biolink predicates account for 58.70% of descriptors, demonstrating the diversity of biomedical relationship types captured in the ontology.
 […]
-tab:top_predicates_distribution_chemprot Predicate Descriptors Percentage (%) substrate 6 20.00 agonist 5 16.67 upregulator 4 13.33 downregulator 4 13.33 regulator 3 10.00 part of 2 6.67 antagonist 2 6.67 modulator 2 6.67 cofactor 2 6.67
-Table 2: Chemprot Dataset Top Predicates by Descriptor Count and Percentage Contribution
+tab:top_predicates_distribution_chemprot Predicate Sample Descriptors Descriptor count (%) substrate Product of, Substrate product of 6 (10.00%) agonist Agonist-activator, Agonist-inhibitor 5 (8.33%) upregulator Activator, Indirect upregulator 4 (6.67%) downregulator Inhibitor, Indirect downregulator 4 (6.67%) regulator Direct regulator, Indirect regulator 3 (5.00%) part of Indicates that the chemical is a structural component or subunit of the protein complex 2 (3.33%) antagonist Chemical substance that binds to and blocks the activation of certain receptors on cells, preventing a biological response 2 (3.33%) modulator Modulator-activator, Modulator-inhibitor 2 (3.33%) cofactor A non-protein chemical compound that is required for an enzyme’s biological activity 2 (3.33%)
+Table 3: ChemProt Dataset.
+All 9 positive relationship predicates ordered by descriptor count.
+ChemProt’s 10th class (NOT/no interaction) is handled via RELATE’s rejection mechanism (NONE option, Stage 3) rather than predicate mapping.
 […]
-tab:top_predicates_distribution_heal Predicate Descriptors Percentage (%) coexists with 178 39.04 related to 96 21.05 located in 40 8.77 part of 34 7.46 temporally related to 30 6.58 has part 21 4.61 precedes 15 3.29 causes 15 3.29 affects 14 3.07 has output 13 2.85
-Table 3: Biolink Dataset Top Predicates by Descriptor Count and Percentage Contribution
-Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal showed the top 10 predicates that account for the majority of descriptors in the dataset.
-substrate and agonist together accounting for over one-third of all descriptors in Table LABEL:tab:top_predicates_distribution_chemprot while Table LABEL:tab:top_predicates_distribution_heal dominated by predicates such as coexists with and related to.
+tab:top_predicates_distribution_heal Predicate Sample Descriptors Descriptor count (%) coexists with holds between two entities that are co-located in the same aggregate object, process, or spatio-temporal region 178 (8.06%) related to holds between two entities 96 (4.39%) located in holds between a material entity and a material entity or site within which it is located (but of which it is not considered a part) 40 (1.81%) part of holds between parts and wholes (material entities or processes) 34 (1.49%) temporally related to holds between two entities with a temporal relationship 30 (1.36%) has part holds between wholes and their parts (material entities or processes) 21 (0.95%) precedes holds between two processes, where one completes before the other begins 15 (0.68%) causes holds between two entities where the occurrence, existence, or activity of one causes the occurrence or generation of the other 15 (0.68%) affects Describes an entity that has an effect on the state or quality of another existing entity 14 (0.63%) has output holds between a process and a continuant, where the continuant is an output of the process 13 (0.63%)
+Table 4: Biolink Dataset. Top 10 of 284 predicates by descriptor count. Percentages calculated from 1,104 total positive descriptors across all 284 Biolink predicates used in this study.